# Get rich quick: exact solutions reveal how unbalanced initializations promote rapid feature learning

### v1

@@ Appendix B Wide and Deep Linear Networks
Vectorising using the identity $vec(\mathbf{ABC})=(\mathbf{C}^{\intercal}\otimes\mathbf{A})vec(\mathbf{B})$ equation 76 becomes
$vec(\dot{\beta})$ $=-vec\left(\eta_{w}I_{d}X^{\intercal}(\underbrace{X\beta-Y}_{P})A^{\intercal}A+\eta_{a}W^{\intercal}WX^{\intercal}(\underbrace{X\beta-Y}_{P})I_{c}\right),$ (77)
$=-(\eta_{w}A^{\intercal}A\otimes I_{c}+\eta_{a}I_{d}\otimes W^{\intercal}W)vec(X^{\intercal}P),$ (78)
$=-\underbrace{\left(\eta_{w}A^{\intercal}A\oplus\eta_{a}W^{\intercal}W\right)}_{M}\mathrm{vec}(X^{\intercal}P).$ (79)
The vectorised form of the network function is given by
$\mathrm{vec}\left(\dot{\beta}\right)=-\underbrace{\left(\eta_{w}A^{\intercal}A\oplus\eta_{a}W^{\intercal}W\right)}_{M}\mathrm{vec}(X^{\intercal}X\beta-X^{\intercal}Y).$ (80)
Interpreting $M(\beta)$ in different limit and architectures
@@
Consider a single hidden neuron $i\in[h]$ of the multi-output model defined by the parameters $w_{i}\in\mathbb{R}^{d}$ and $a_{i}\in\mathbb{R}^{c}$. Let $\beta_{i}=w_{i}a_{i}^{\intercal}$ be the $\mathbb{R}^{d\times c}$ matrix representing the contribution of this hidden neuron to the input-output map of the network. As in the previous section, we will consider the two gram matrices $\beta_{i}^{\intercal}\beta_{i}\in\mathbb{R}^{c\times c}$ and $\beta_{i}\beta_{i}^{\intercal}\in\mathbb{R}^{d\times d}$,
$\beta_{i}^{\intercal}\beta_{i}=\|w_{i}\|^{2}a_{i}a_{i}^{\intercal},\qquad\beta_{i}\beta_{i}^{\intercal}=\|a_{i}\|^{2}w_{i}w_{i}^{\intercal}.$ (81)
Notice that we can express $\|\beta_{i}\|_{F}^{2}$ as
$\|\beta_{i}\|_{F}^{2}=\mathrm{Tr}(\beta_{i}^{\intercal}\beta_{i})=\mathrm{Tr}(\beta_{i}\beta_{i}^{\intercal})=\|a_{i}\|^{2}\|w_{i}\|^{2}$ (82)
@@
At each hidden neuron we have the conserved quantity [^8] As long as $c>1$, then the surface of this $d+c$ hyperboloid is always connected, however its topology will depend on the relationship between $d$ and $c$. $\eta_{w}\|a_{i}\|^{2}-\eta_{a}\|w_{i}\|^{2}=\delta_{i}$ where $\delta_{i}\in\mathbb{R}$. Using this quantity we can invert the expression for $\|\beta_{i}\|_{F}^{2}$ to get
$\|a_{i}\|^{2}$ $=\frac{\sqrt{\delta_{i}^{2}+\eta_{a}\eta_{w}4\|\beta_{i}\|_{F}^{2}}+\delta_{i}}{2},$ (83)
$\|w_{i}\|^{2}$ $=\frac{\sqrt{\delta_{i}^{2}+\eta_{a}\eta_{w}4\|\beta_{i}\|_{F}^{2}}-\delta_{i}}{2}.$ (84)
@@
When $\|\beta_{i}\|_{F}^{2}>0$, we can use these expressions to solve for the outer products $a_{i}a_{i}^{\intercal}$ and $w_{i}w_{i}^{\intercal}$ entirely in terms of $\beta_{i}$,
@@
$a_{i}a_{i}^{\intercal}$ $=\frac{\sqrt{\delta_{i}^{2}+\eta_{a}\eta_{w}4\|\beta_{i}\|_{F}^{2}}+\delta_{i}}{2}\frac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}},$ (85)
$w_{i}w_{i}^{\intercal}$ $=\frac{\sqrt{\delta_{i}^{2}+\eta_{a}\eta_{w}4\|\beta_{i}\|_{F}^{2}}-\delta_{i}}{2}\frac{\beta_{i}\beta_{i}^{\intercal}}{\|\beta_{i}\|_{F}^{2}}.$ (86)
Without making any assumptions on the initialization (such as the isotropic initialization) we can express the NTK in terms of the $\beta_{i}$ and consider the effect the vector of conserved quantities $\delta\in\mathbb{R}^{h}$ has on the dynamics.
@@
Assuming $\beta_{i}\neq 0$ for all $i\in[h]$ and let $\kappa_{i}=\sqrt{\delta_{i}^{2}+4\eta_{a}\eta_{w}\|\beta_{i}\|_{F}^{2}}$, then the matrix $M$ can be expressed as the sum $M=\sum_{i=1}^{h}M_{i}$ over hidden neurons where $M_{i}$ is defined as,
$M_{i}=\left(\frac{\kappa_{i}+\delta_{i}}{2}\right)\frac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}\oplus\left(\frac{\kappa_{i}-\delta_{i}}{2}\right)\frac{\beta_{i}\beta_{i}^{\intercal}}{\|\beta_{i}\|_{F}^{2}}.$ (87)
We show how to express $M$ in terms of the matrices $\beta_{i}=w_{i}a_{i}^{\intercal}\in\mathbb{R}^{d\times c}$, which represent the contribution to the input-output map of a single hidden neuron $i\in[h]$ of the network with parameters $w_{i}\in\mathbb{R}^{d}$, $a_{i}\in\mathbb{R}^{c}$, and conserved quantity $\delta_{i}=\Delta_{ii}$.
B.1.1 Funnel Networks
We consider funnel networks, which narrow from input to output ($d>h\geq c$), and inverted-funnel networks, which expand from input to output ($d\leq h<c$).
As $\delta_{i}\to\infty$, $M\to\sum_{i=1}^{h}|\delta_{i}|\tfrac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}\otimes\mathbf{I}_{d}$.
@@ B.1.1 Funnel Networks
Consider the rank-one matrices $\frac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}$ in the space $\mathbb{R}^{c\times c}$. The rank of M is bounded by
$\operatorname{rank}\left(M\right)\leq\begin{cases}d\times h&\textit{if }h<c\textit{, }\text{Low-rank}\\ d\times c&\textit{if }h\geq c\textit{. }\end{cases}$ (88)
@@
In this limit, the rank of M is given by
$d\times\operatorname{rank}(\sum_{i=1}^{h}\tfrac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}})\leq d\times\sum_{i=1}^{h}\operatorname{rank}(\tfrac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}).$ (89)
It follows that
$\sum_{i=1}^{h}\operatorname{rank}\left(\frac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}\right)\leq\begin{cases}h&\textit{if }h<c.\\ c&\textit{if }h\geq c.\end{cases}$ (90)
as the rank of the sum $\sum_{i=1}^{h}\frac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}$ is also at most $\operatorname{rank}c$. Therefore,
$\operatorname{rank}\left(M\right)\leq\begin{cases}d\times h&\textit{if }h<c.\\ d\times c&\textit{if }h\geq c.\end{cases}$ (91)
According to lemma B.2
  • If $h<c$, $\operatorname{rank}(M)$ is bounded by $d\times h$ and remains below $d\times c$, categorizing it as a low-rank matrix. As a result, the solution $\beta_{*}^{\intercal}\beta_{*}$ might be in the null space of $M$. The network may enter either the lazy or a lazy followed by rich regime, depending on the relationship between $\beta_{0}$ and $\beta_{*}$. If $\beta_{0}\propto\beta_{*}$ the network will enter the lazy regime.
  • Assuming the $\tfrac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}$ terms are linearly independent, the matrix $\operatorname{rank}(M)$ achieves full $\operatorname{rank}$ and spans the solution space. The network can learn the task by only changing their norm while keeping their direction and the NTK matrix fixed. Thus, funnel networks defined by ($d>h\geq c$) will transition into the lazy regime in this limit.
A similar assertion applies as $\delta_{i}\to-\infty$. In this limit, $M\to$ $\mathbf{I}_{c}\otimes\sum_{i=1}^{h}|\delta_{i}|\frac{\beta_{i}\beta_{i}^{\intercal}}{\|\beta_{i}\|_{F}^{2}}$ with the $\operatorname{rank}(M)$ being constrained by the relationship between the number of hidden layers $h$ and the input layer dimensions $d$.
  • If $h<d$, the matrix $M$ is low $\operatorname{rank}$ and bounded by $c\times h$. These networks may enter either the lazy or a lazy followed by rich regime, depending on the relationship between $\beta_{0}$ and $\beta_{*}$.
  • When $h\geq d$, $M$ is full $\operatorname{rank}$ if $\tfrac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}$ are linearly independent. Except for a measure zero set of initializations, inverted funnel networks always enter the lazy regime in this limit.
When $\delta_{i}=0$, $M=\sqrt{\eta_{a}\eta_{w}}|\beta_{i}|\sum_{i=1}^{h}\tfrac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}\oplus\sum_{i=1}^{h}\tfrac{\beta_{i}\beta_{i}^{\intercal}}{\|\beta_{i}\|_{F}^{2}}$ all networks transition into the rich regime. Employing a similar rationale as before, assuming that all terms $\frac{\beta_{i}^{\intercal}\beta_{i}}{|\beta_{i}|_{F}^{2}}$ and $\frac{\beta_{i}\beta_{i}^{\intercal}}{|\beta_{i}|_{F}^{2}}$ are respectively linearly separable, then one term of $M$ will be low-rank, while the other will be full-rank, contingent on the relationship between $c$, $d$, and $h$. Consequently, the dynamics of the network balance between low-rank and full-rank elements, leading to changes in both the magnitude and direction of the Neural Tangent Kernel (NTK) during training, a hallmark of the rich regime.
B.1.2 Single-Neuron
For funnel network with a single hidden neuron ($h=c=1$), we recover equation 4 from equation 7.We extend this analysis to inverted-funnel network with a single hidden neuron ($h=d=1$). Assuming $h=c=1$, the rank one matrix $\frac{\beta_{i}^{\intercal}\beta_{i}}{\|\beta_{i}\|_{F}^{2}}=1$. Therefore, equation 7 becomes
$M$ $=\left(\frac{\kappa+\delta}{2}\right)\oplus\left(\frac{\kappa-\delta}{2}\right)\frac{\beta\beta^{\intercal}}{\|\beta\|_{F}^{2}}.$ (92)
$=\frac{\kappa+\delta}{2}I_{d}+\frac{\kappa-\delta}{2}\frac{\beta\beta^{\intercal}}{\|\beta\|^{2}}.$ (93)
where $\kappa=\sqrt{\delta^{2}+\eta_{a}\eta_{w}4\|\beta\|^{2}}$. We recover the funnel network equation 4 for a single-neuron. In the main text, we analyze how the expression for $M(\beta)$ simplifies when $\delta$ approaches $\infty$- ∞, 0, and $\infty$. This analysis helps us develop a deeper understanding of $M_{\delta}(\beta)$.
We now turn to single neuron inverted funnel network where $h=d=1$, the rank one matrices $\frac{\beta_{i}\beta_{i}^{\intercal}}{\|\beta_{i}\|_{F}^{2}}=1$. Therefore, equation 7 becomes
$M$ $=\left(\frac{\kappa+\delta}{2}\right)\frac{\beta^{\intercal}\beta}{\|\beta\|_{F}^{2}}\oplus\left(\frac{\kappa-\delta}{2}\right)$ (94)
$=\frac{\kappa+\delta}{2}\frac{\beta^{\intercal}\beta}{\|\beta\|_{F}^{2}}+\frac{\kappa-\delta}{2}I_{c}.$ (95)
From our expression for $M(\beta)$ we will consider how it simplifies when $\delta\to-\infty,0,\infty$.
$M\to\begin{cases}\delta I_{c}&\delta\to-\infty.\\ \sqrt{\eta_{a}\eta_{w}}\|\beta\|\left(\frac{\beta^{\intercal}\beta}{\|\beta\|^{2}}+I_{c}\right)&\delta=0.\\ |\delta|\frac{\beta^{\intercal}\beta}{\|\beta\|^{2}}&\delta\to\infty.\end{cases}$ (96)
As anticipated, the single-neuron inverted funnel network also transitions into the Rich regime when $\delta=0$. Under this condition, $M=\sqrt{\eta_{a}\eta_{w}}\|\beta\|(\frac{\beta^{\intercal}\beta}{\|\beta\|^{2}}+I_{c})$, where the initial term denotes a projection matrix. Here the dynamics balance between following the lazy trajectory and attempting to fit the task by only changing in norm. As a result the NTK changes in both magnitude and direction through training, confirming the dynamics are rich.
Additionally, the single-neuron inverted funnel network follows analogous regime transitions to the single-neuron funnel network, albeit in opposite directions. As $\delta\to-\infty$, $M\to\delta I_{c}$, the single-neuron inverted funnel network network enters the Lazy regime. In this regime, the dynamics of $\beta$ converge to the trajectory of linear regression trained by gradient flow and along this trajectory the NTK matrix remains constant.
Conversely, as $\delta\to+\infty$, $M\to|\delta|\frac{\beta^{\intercal}\beta}{\|\beta\|^{2}}$, transitions into the Lazy-to-Rich regime for the network. Here the dynamics of $\beta$ are constrained to learn the task by only changing their norm while keeping the direction and the NTK matrix fixed – an initial lazy phase. However, if $\beta$ must change direction to fit the task, and assuming finite $\delta$ or that $t\to\infty$ faster than $\delta\to-\infty$, then at some point there will be a slow alignment of $\beta$ to $\beta_{*}$. In this second phase the NTK matrix will change, confirming the dynamics are lazy-to-rich.
In summary, networks with scalar outputs enter the lazy regime as $\delta\rightarrow\infty$, while networks with scalar inputs enter the lazy regime as $\delta\rightarrow-\infty$. Conversely, both types of networks enter the "active" regime in opposite directions. Finally, both cases of scalar output and scalar input enter the rich regime when $\delta\to 0$
B.1.3 Multi-Output
[…]
We consider the isotropic initialization defined as $\Delta=\delta I_{h}$ in this section. The conserved quantity becomes $AA^{\intercal}-WW^{\intercal}=\delta I_{h}$.
Lemma B.3.
Assuming an initialization where $\|\beta(0)\|_{F}>0$ and $\Delta=\delta I_{h}$, then the dynamics of the network can be expressed as $vec(\beta)=-Mvec(X^{\intercal}P)$, where $M$ is defined as
$M$ $=\left(\sqrt{\eta_{a}\eta_{w}\beta^{\intercal}\beta+\frac{\delta^{2}}{4}I}\otimes I\right)+\left(I\otimes\sqrt{\eta_{a}\eta_{w}\beta\beta^{\intercal}+\frac{\delta^{2}}{4}I}\right).$ (97)
@@ B.1.3 Multi-Output
With some algebras, one can show the following relations.
$W^{\intercal}W$ $=\frac{1}{\eta_{a}}\left(-\frac{\delta}{2}I+\sqrt{\eta_{a}\eta_{w}\beta\beta^{\intercal}+\frac{\delta^{2}}{4}I}\right)$ (98)
$A(t)^{\intercal}A(t)$ $=\frac{1}{\eta_{w}}\left(\frac{\delta}{2}I+\sqrt{\eta_{a}\eta_{w}\beta^{\intercal}\beta+\frac{\delta^{2}}{4}I}\right).$ (99)
@@
[…]
@@
$vec(\dot{\beta})$ $=-\left[\underbrace{\left(\sqrt{\eta_{a}\eta_{w}\beta^{\intercal}\beta+\frac{\delta^{2}}{4}I}\otimes I\right)+\left(I\otimes\sqrt{\eta_{a}\eta_{w}\beta\beta^{\intercal}+\frac{\delta^{2}}{4}I}\right)}_{M}\right]vec(X^{\intercal}P).$ (101)

### v2

@@ B.1.2 Function space dynamics
Vectorizing using the identity $\mathrm{vec}(ABC)=(C^{\intercal}\otimes A)\mathrm{vec}(B)$ equation 85 becomes
$\mathrm{vec}\left(\dot{\beta}\right)$ $=-\mathrm{vec}\left(\eta_{w}\mathbf{I}_{d}X^{\intercal}(X\beta-Y)A^{\intercal}A+\eta_{a}W^{\intercal}WX^{\intercal}(X\beta-Y)\mathbf{I}_{c}\right),$ (86)
$=-(\eta_{w}A^{\intercal}A\otimes\mathbf{I}_{d}+\eta_{a}\mathbf{I}_{c}\otimes W^{\intercal}W)\mathrm{vec}(X^{\intercal}X\beta-X^{\intercal}Y),$ (87)
$=-\underbrace{\left(\eta_{w}A^{\intercal}A\oplus\eta_{a}W^{\intercal}W\right)}_{M}\mathrm{vec}(X^{\intercal}X\beta-X^{\intercal}Y).$ (88)
@@
We first prove Theorem 4.1. Consider a single hidden neuron $k\in[h]$ of the multi-output model defined by the parameters $w_{k}\in\mathbb{R}^{d}$ and $a_{k}\in\mathbb{R}^{c}$. Let $\beta_{k}=w_{k}a_{k}^{\intercal}$ be the $\mathbb{R}^{d\times c}$ matrix representing the contribution of this hidden neuron to the input-output map of the network $\beta=\sum_{k=1}^{h}\beta_{k}$. Consider the two gram matrices $\beta_{k}^{\intercal}\beta_{k}\in\mathbb{R}^{c\times c}$ and $\beta_{k}\beta_{k}^{\intercal}\in\mathbb{R}^{d\times d}$,
$\beta_{k}^{\intercal}\beta_{k}=\|w_{k}\|^{2}a_{k}a_{k}^{\intercal},\qquad\beta_{k}\beta_{k}^{\intercal}=\|a_{k}\|^{2}w_{k}w_{k}^{\intercal}.$ (89)
Notice that we can express $\|\beta_{k}\|_{F}^{2}$ as
$\|\beta_{k}\|_{F}^{2}=\mathrm{Tr}(\beta_{k}^{\intercal}\beta_{k})=\mathrm{Tr}(\beta_{k}\beta_{k}^{\intercal})=\|a_{k}\|^{2}\|w_{k}\|^{2}$ (90)
@@
At each hidden neuron we have the conserved quantity [^8] As long as $c>1$, then the surface of this $d+c$ hyperboloid is always connected, however its topology will depend on the relationship between $d$ and $c$. $\eta_{w}\|a_{k}\|^{2}-\eta_{a}\|w_{k}\|^{2}=\delta_{k}$ where $\delta_{k}\in\mathbb{R}$. Using this quantity we can invert the expression for $\|\beta_{k}\|_{F}^{2}$ to get
$\|a_{k}\|^{2}$ $=\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\|\beta_{k}\|_{F}^{2}}+\delta_{k}}{2\eta_{w}},$ (91)
$\|w_{k}\|^{2}$ $=\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\|\beta_{k}\|_{F}^{2}}-\delta_{k}}{2\eta_{a}}.$ (92)
@@
When $\|\beta_{k}\|_{F}^{2}>0$, we can use these expressions to solve for the outer products $a_{k}a_{k}^{\intercal}$ and $w_{k}w_{k}^{\intercal}$ in terms of $\beta_{k}$ and $\delta_{k}$,
$a_{k}a_{k}^{\intercal}$ $=\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\|\beta_{k}\|_{F}^{2}}+\delta_{k}}{2\eta_{w}}\frac{\beta_{k}^{\intercal}\beta_{k}}{\|\beta_{k}\|_{F}^{2}},$ (93)
$w_{k}w_{k}^{\intercal}$ $=\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\|\beta_{k}\|_{F}^{2}}-\delta_{k}}{2\eta_{a}}\frac{\beta_{k}\beta_{k}^{\intercal}}{\|\beta_{k}\|_{F}^{2}}.$ (94)
By substituting these expressions into the decompositions $A^{\intercal}A=\sum_{k=1}^{h}a_{k}a_{k}^{\intercal}$ and $W^{\intercal}W=\sum_{k=1}^{h}w_{k}w_{k}^{\intercal}$, we derive the representation for $M$ presented in Theorem 4.1: $M=\sum_{k=1}^{h}M_{k}$ where
@@
$M_{k}=\left(\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\|\beta_{k}\|_{F}^{2}}+\delta_{k}}{2}\right)\frac{\beta_{k}^{\intercal}\beta_{k}}{\|\beta_{k}\|_{F}^{2}}\oplus\left(\frac{\sqrt{\delta_{k}^{2}+4\eta_{a}\eta_{w}\|\beta_{k}\|_{F}^{2}}-\delta_{k}}{2}\right)\frac{\beta_{k}\beta_{k}^{\intercal}}{\|\beta_{k}\|_{F}^{2}}.$ (95)
B.1.4 Understanding $M$ when there is a single-neuron $h=1$
When there is a single-hidden neuron $h=\min(d,c)=1$, the expression for $M$ presented in Theorem 4.1 simplifies allowing us to precisely understand the influence of $\delta$ on the learning regime. When $h=c=1$, then $\frac{\beta^{\intercal}\beta}{\|\beta\|_{F}^{2}}=1$. Therefore, Eq. 7 simplifies to
$M=\frac{\sqrt{\delta^{2}+\eta_{a}\eta_{w}4\|\beta\|^{2}}+\delta}{2}\mathbf{I}_{d}+\frac{\sqrt{\delta^{2}+\eta_{a}\eta_{w}4\|\beta\|^{2}}-\delta}{2}\frac{\beta\beta^{\intercal}}{\|\beta\|^{2}},$ (96)
and we recover Eq. 4 presented in Section 3. When $h=d=1$, then $\frac{\beta\beta^{\intercal}}{\|\beta\|_{F}^{2}}=1$ and thus Eq. 7 simplifies to,
$M=\frac{\sqrt{\delta^{2}+\eta_{a}\eta_{w}4\|\beta\|^{2}}+\delta}{2}\frac{\beta^{\intercal}\beta}{\|\beta\|^{2}}+\frac{\sqrt{\delta^{2}+\eta_{a}\eta_{w}4\|\beta\|^{2}}-\delta}{2}\mathbf{I}_{c}.$ (97)
In both settings, $M$ is the weighted sum of the identity matrix and a rank-one projection matrix. While these equations are strikingly similar there is an interesting distinction that arises in the limits of $\delta$. As $\delta\to\infty$, then the first expression for $M$ becomes proportional to $\mathbf{I}_{d}$, while the second expression for $M$ becomes proportional to the rank-1 projection $\frac{\beta^{\intercal}\beta}{\|\beta\|^{2}}$. Conversely, as $\delta\to-\infty$, then the first expression for $M$ becomes proportional to the rank-1 projection $\frac{\beta\beta^{\intercal}}{\|\beta\|^{2}}$, while the second expression for $M$ becomes proportional to $\mathbf{I}_{c}$. When $h=d=c=1$, then $M=\sqrt{\delta^{2}+\eta_{a}\eta_{w}4\|\beta\|^{2}}$ and thus in both limits of $\delta\to\pm\infty$, $M$ becomes a constant independent of $\beta$. In all settings, when $\delta=0$, $M$ depends on $\beta$. In other words, the influence of $\delta$ on whether the dynamics are lazy, rich, or delayed rich, crucially depends on the relative sizes of dimensions $d$, $h$, and $c$.
B.1.5 Interpreting $M$ in different limits and architectures
We now seek to more generally understand the influence of the conserved quantities $\delta_{i}$ and the relative sizes of dimensions $d$, $h$ and $c$ on the learning regime. For a matrix $A\in\mathbb{R}^{d\times c}$, let $\mathrm{Row}(A)\subseteq\mathbb{R}^{c}$ and $\mathrm{Col}(A)\subseteq\mathbb{R}^{d}$ denote the row and column space of $A$ respectively.
Theorem B.1.
The dynamics are in the lazy regime, for all $t\geq 0$, if $\delta_{k}\to\infty$ for all $k\in[h]$ and there exists a least squares solution $\beta_{*}\in\mathbb{R}^{d\times c}$ such that
$\mathrm{Row}(\beta_{*})\subseteq\mathrm{Span}\left(\bigcup_{k=1}^{h}\mathrm{Row}\left(\beta_{k}(0)\right)\right),$ (98)
or $\delta_{k}\to-\infty$ for all $k\in[h]$ and there exists a solution such that
$\mathrm{Col}(\beta_{*})\subseteq\mathrm{Span}\left(\bigcup_{k=1}^{h}\mathrm{Col}\left(\beta_{k}(0)\right)\right).$ (99)
Proof.
As $\delta_{k}\to\infty$, $M_{k}\to|\delta_{k}|\tfrac{\beta_{k}^{\intercal}\beta_{k}}{\|\beta_{k}\|_{F}^{2}}\otimes\mathbf{I}_{d}$, implying $\dot{\beta_{k}}=-|\delta_{k}|\frac{\partial\mathcal{L}}{\partial\beta}\left(\tfrac{\beta_{k}^{\intercal}\beta_{k}}{\|\beta_{k}\|_{F}^{2}}\right)$. Notice that $\left(\tfrac{\beta_{k}^{\intercal}\beta_{k}}{\|\beta_{k}\|_{F}^{2}}\right)$ is the unique orthogonal projection matrix onto the one-dimensional row space of $\beta_{k}$. Thus, the dynamics of each $\beta_{k}$ follow a projected gradient descent in their row space. As a result, $M_{k}$ will not change and thus the NTK will be static. By assumption there exists a least squares solution $\beta_{*}$ such that the rows of $\beta_{*}$ are in the span of the rows of $\beta_{k}$. Thus, a solution will be reached as $t\to\infty$, while the $M_{k}$ remain static.
As $\delta_{k}\to-\infty$ for all $k\in[h]$, $M_{k}\to\mathbf{I}_{c}\otimes|\delta_{k}|\tfrac{\beta_{k}\beta_{k}^{\intercal}}{\|\beta_{k}\|_{F}^{2}}$, and an analogous argument can be made. ∎
Note that the assumptions in Theorem B.1 can be more intuitively expressed in terms of the parameter space $(W,A)$. Except in highly degenerate cases, the assumption $\mathrm{Row}(\beta_{*})\subseteq\mathrm{Span}\left(\bigcup_{k=1}^{h}\mathrm{Row}\left(\beta_{k}(0)\right)\right)$ is equivalent to the existence of a $\beta_{*}$ whose rows lie in the span of $\{a_{k}(0)\}_{k=1}^{h}$, or, equivalently, to the existence of a matrix $W$ such that $\beta_{*}=W^{\intercal}A(0)$. Similarly, the condition $\mathrm{Col}(\beta_{*})\subseteq\mathrm{Span}\left(\bigcup_{k=1}^{h}\mathrm{Col}\left(\beta_{k}(0)\right)\right)$ is in most cases equivalent to the existence of a matrix $A$ such that $\beta_{*}=W(0)^{\intercal}A$.
A direct consequence of Theorem B.1 is that networks which narrow from input to output ($d>c$) must enter the lazy regime with probability 1 as all $\delta_{k}\rightarrow\infty$ whenever $h\geq c$ and assuming independent initializations for all $\beta_{k}$. In this case, the rows of $\{\beta_{1},\dots,\beta_{h}\}$ span all of $\mathbb{R}^{c}$ and thus the condition on the least squares solution is trivially true. By the same logic, networks which expand from input to output ($d<c$) do so as all $\delta_{k}\to-\infty$ whenever $h\geq d$ and assuming independent initializations for all $\beta_{k}$. Additionally, when $h\geq\max(d,c)$ and assuming independent initializations for all $\beta_{k}$, then all networks enter the lazy regime as either all $\delta_{k}\to\infty$ or all $\delta_{k}\to-\infty$.
Another interesting implication of Theorem B.1, is that if there does not exist a least squares solution $\beta_{*}$ with rows in the span of the rows of $\{\beta_{1},\dots,\beta_{h}\}$, then the network will enter a delayed rich regime as all $\delta_{k}\to\infty$, where the magnitude of the $\delta_{k}$ will determine the delay. In this setting, the network is initially lazy, attempting to fit the solution within the row space of the $\beta_{k}$, but eventually the direction of the rows must change in order to fit the problem, leading to a rich phase. A similar statement involving the columns of $\beta_{*}$ is true as all $\delta_{k}\to-\infty$.
B.1.6 Simplifying $M$ through assumptions on $\Delta$
We now consider how introducing structures on $\Delta$ can lead to simpler expressions for $M$. A natural assumption to consider is the following:
Assumption B.2 (Isotropic initialization).
Let $A\in\mathbb{R}^{h\times c}$ and $W\in\mathbb{R}^{h\times d}$ be initialized such that $\Delta=\eta_{w}A(0)A(0)^{\intercal}-\eta_{a}W(0)W(0)^{\intercal}=\delta\mathbf{I}_{h}$.
In square networks, where the dimensions of the input, hidden, and output layers coincide ($d=h=c$), and the weights are initialized as $A\sim\mathcal{N}(0,\sigma_{a}^{2}/c)$ and $W\sim\mathcal{N}(0,\sigma_{w}^{2}/d)$, this assumption is naturally satisfied with $\delta=\sigma_{a}^{2}-\sigma_{w}^{2}$ as the dimension $h\to\infty$. However, a limitation of this assumption is that for general $\delta$ it requires $h\leq\min(d,c)$. Specifically, when $\delta>0$, the isotropic initialization requires that $A(0)A(0)^{\intercal}\succ 0$, which implies $h\leq c$. Similarly, when $\delta<0$, the isotropic initialization requires that $W(0)W(0)^{\intercal}\succ 0$, which implies $h\leq d$. Now we prove two important implications of the isotropic initialization assumption.
@@ B.1.6 Simplifying $M$ through assumptions on $\Delta$
Let $\Delta=\delta\mathbf{I}_{h}$. If either $\delta\geq 0$ or $\delta<0$ and $h\geq d$, we have that
$W^{\intercal}W=\frac{1}{\eta_{a}}\left(-\frac{\delta}{2}\mathbf{I}_{d}+\sqrt{\eta_{a}\eta_{w}\beta\beta^{\intercal}+\frac{\delta^{2}}{4}\mathbf{I}_{d}}\right).$ (100)
Proof.
The quantity $\eta_{w}AA^{\intercal}-\eta_{a}WW^{\intercal}=\delta\mathbf{I}_{h}$ is conserved in gradient flow. Multiplying on the left by $W^{\intercal}$ and on the right by $W$ we have that
$\eta_{a}(W^{\intercal}W)^{2}+\delta W^{\intercal}W=\eta_{w}\beta\beta^{\intercal}.$ (101)
Completing the square by adding $\frac{\delta^{2}}{4\eta_{a}}\mathbf{I}_{d}$ to both sides and dividing by $\eta_{a}$ we get the equality,
$\left(W^{\intercal}W+\frac{\delta}{2\eta_{a}}\mathbf{I}_{d}\right)^{2}=\frac{\delta^{2}}{4\eta_{a}^{2}}\mathbf{I}_{d}+\frac{\eta_{w}}{\eta_{a}}\beta\beta^{\intercal}$ (102)
For $\delta\geq 0$, $W^{\intercal}W+\frac{\delta}{2\eta_{a}}\mathbf{I}_{d}\succeq 0$. For $\delta<0$, then we know from the conserved quantity that $WW^{\intercal}+\frac{\delta}{2\eta_{a}}\mathbf{I}_{h}=\frac{\eta_{w}}{\eta_{a}}AA^{\intercal}-\frac{\delta}{2\eta_{a}}\mathbf{I}_{h}\succ 0$, which implies when $h\geq d$ that $W^{\intercal}W+\frac{\delta}{2\eta_{a}}\mathbf{I}_{d}\succ 0$. As a result, we can take the principal square root of each side,
$W^{\intercal}W+\frac{\delta}{2\eta_{a}}\mathbf{I}_{d}=\sqrt{\frac{\delta^{2}}{4\eta_{a}^{2}}\mathbf{I}_{d}+\frac{\eta_{w}}{\eta_{a}}\beta\beta^{\intercal}},$ (103)
which rearranged gives the final result. ∎
@@
Let $\Delta=\delta\mathbf{I}_{h}$. If either $\delta\leq 0$ or $\delta>0$ and $h\geq c$, we have that
$A^{\intercal}A=\frac{1}{\eta_{w}}\left(\frac{\delta}{2}\mathbf{I}_{c}+\sqrt{\eta_{a}\eta_{w}\beta^{\intercal}\beta+\frac{\delta^{2}}{4}\mathbf{I}_{c}}\right).$ (104)
@@
The proof is analogous to the proof of Lemma B.3.
From Lemma B.3 and Lemma B.4 we can prove Theorem 4.2, as shown below.
@@
Plugging in expressions for $W^{\intercal}W$ from Lemma B.3 and $A^{\intercal}A$ from Lemma B.4 we can directly write,
@@
$M$ $=\left(\frac{\delta}{2}\mathbf{I}_{c}+\sqrt{\eta_{a}\eta_{w}\beta^{\intercal}\beta+\frac{\delta^{2}}{4}\mathbf{I}_{c}}\right)\oplus\left(-\frac{\delta}{2}\mathbf{I}_{d}+\sqrt{\eta_{a}\eta_{w}\beta\beta^{\intercal}+\frac{\delta^{2}}{4}\mathbf{I}_{d}}\right)$ (106)
$=\left(\sqrt{\eta_{a}\eta_{w}\beta^{\intercal}\beta+\frac{\delta^{2}}{4}\mathbf{I}_{c}}\otimes\mathbf{I}_{d}\right)+\left(\mathbf{I}_{c}\otimes\sqrt{\eta_{a}\eta_{w}\beta\beta^{\intercal}+\frac{\delta^{2}}{4}\mathbf{I}_{d}}\right)$ (107)